OpenStudy (anonymous):

Write the quadratic equation in standard form 2(x − 3)2 = 1

OpenStudy (anonymous):

2(x-3)^2=1 2[x^2 - 6x + 9] = 1 2x^2 - 12x + 18 = 1 2x^2 - 12x + 17 = 0

OpenStudy (anonymous):

that 2 at the end of 2(x-3)2 is supposed to be 2(x-3)^2 right?

OpenStudy (anonymous):

yes jpomer

OpenStudy (anonymous):

alright then thats your answer

OpenStudy (anonymous):

you get it?

OpenStudy (anonymous):

no